# Investigating Disordered Granular Matter via Ordered Geometric Fragmentation

### v1

@@ III.5 Remarks on paired configurations and “phase” interpretation
As demonstrated below, the pairing structure is closely related to the symmetry properties of the tower configurations.
@@
Since $i$ is a non-negative whole number, this condition requires careful interpretation. If $n$ is even, Eq. (13) yields an integer value, and the pair (i,n+2i)(i,n+2-i) reduces to $(n/2+1,n/2+1)$. In this case, two conjugate towers coincide and form a single self-conjugate tower. If $n$ is odd, such a reduction does not occur, and a self-conjugate tower never exists.
Conjugate towers are characterized by a rearrangement of identical building blocks. When these towers coincide in the even-$n$ case, the resulting self-conjugate tower inherits properties of both configurations. It can therefore be inferred that the self-conjugate tower may be constructed from the same building blocks in two distinct orientations, implying the presence of an internal symmetry, in agreement with the behaviour already observed in the example $n=2$ discussed in the subsection Case $n=2(l=4a)$.
These paired configurations provide a simple geometric analogue of distinct phases. While they are composed of identical building blocks, they differ by a global rearrangement that modifies the enclosed cavity volume. Rearrangement-induced transitions between conjugate towers thus correspond to transitions between geometrically distinct configurations with different effective densities.
The geometric transitions observed in the present model admit a natural interpretation in the language of phase transitions. The relative volume $R_{n}^{\,i}$ plays the role of a macroscopic order parameter distinguishing tower configurations. Transitions between conjugate towers involve discontinuous changes in $R_{n}^{\,i}$ and may therefore be regarded as first-order-like. In contrast, when $n$ is even, the middle tower is self-conjugate: rearrangement of the building blocks alters their internal orientation without changing the overall geometry or volume. This transition is associated with a discrete rotational symmetry and corresponds to a change in internal symmetry rather than in macroscopic structure. In this sense, the self-conjugate tower represents a geometric analogue of a second-order phase transition.
@@
IV Physical interpretation and connections to granular matter
To connect the present model with real granular systems, Eq. (10) is employed to evaluate the relative volume accessible to grains in experiment, while Eq. (12) is employed to identify the geometric phase transitions predicted by the model. Both expressions are formulated in terms of the discrete parameters $n$ and $i$, which must therefore be re-expressed in terms of experimentally measurable grain characteristics.
@@ IV.1 Remarks on experimental observation of the relative volume
A granular sample composed of identical grains with square cross-section of side length $a$ is considered. It is further assumed that these grains may be regarded as having been produced by sequential slicing of a hypothetical elongated parent object, in accordance with the fragmentation rule introduced earlier. By comparing the total mass of the granular sample with the mass of an individual grain, the total number of grains in the sample, denoted by $N_{g}$, is determined. Let $l_{g}$ denote the length of an individual grain. The length $l$ of the hypothetical parent object is then given by
@@
Our next goal is to identify the fragmentation stage of the hypothetical parental object at which the size of a model grain coincides with that of the grains used in the experiment. At fragmentation stage $i$, the grain length $l_{i}$ (given by Eq. (6)) must match the experimental grain length $l_{g}$, so that
@@
As a consistency check, it is noted that for a single grain ($N_{g}=1$), one obtains $i=0$, corresponding to the absence of fragmentation. Substituting Eqs. (15) and (16) into Eq. (10) yields
@@
Although the relative volume $R$ is expressed in Eq. (17) as a single function of the geometric ratio $l_{g}/a$, this quantity should be understood as a characteristic upper bound rather than as a uniquely reproducible experimental outcome. Under nominally identical preparation conditions, repeated realizations of the system are expected to yield a distribution of measured relative volumes clustered below this bound. For a granular assembly composed of grains with square cross section, the relative volume may therefore be determined experimentally by measuring only the grain length $l_{g}$ and the side length $a$ of the cross section. Upon successive fragmentation, these measurements may be repeated until the terminal stage is reached, at which $l_{g}=a$. In this limit, the relative volume approaches the value $5/4$, in agreement with Eq. (11). Exact quantitative agreement with the theoretical curve is not expected under experimental conditions. Nevertheless, the measured values are expected to follow the predicted qualitative trend and to remain bounded above by the maximal-volume curve provided by the model. Assuming square cross sections, these considerations may be summarised as follows.
@@
If a grain has length $l_{g}$ and a square cross section with side length $a$, the upper bound of the relative volume is determined solely by the ratio $l_{g}/a$.
@@
The upper bound of the relative volume decreases monotonically as the grain length $l_{g}$ decreases. Its absolute minimum is attained when $l_{g}=a$, where the upper bound equals $5/4$, a universal value referred to as Liza’s limit or Liza’s number. The upper bound always exceeds unity, corresponding to the volume obtained by packing all grains adjacently without voids.
Equation (17) may be tested against existing experimental data. In Ref. [9], experiments were conducted on monodisperse nylon $6/6$ cylindrical rods with diameter $1.8\,\mathrm{mm}$ and length $7.0\,\mathrm{mm}$. Approximately $7200$ rods were poured into a vertical glass tube of height $1\,\mathrm{m}$ and inner diameter $1.90\,\mathrm{cm}$. Although the present model is formulated for grains with square cross section, a reasonable estimate is expected to be obtained for cylindrical rods of comparable transverse dimensions. The rods are therefore approximated as grains with square cross section of side length a=1.8a=1.8mm and grain length $l_{g}=7.0\,\mathrm{mm}$. Substitution into Eq. (17) then yields
@@
Experimentally, the quantity reported is the packing factor $\rho$ (also commonly denoted $\phi$), defined as the fraction of the available volume occupied by the particles. When the longest grain dimension is small compared to the characteristic dimensions of the container — in the present case, the rod length relative to the tube diameter — the packing factor may be approximated as $\rho\simeq R^{-1}$. This estimate gives $\rho\simeq 0.51$, which lies within the experimentally observed range $\rho=0.49$$0.55$ reported in Ref. [9].
IV.2 Remarks on experimental observation of the “phase transition”
@@
An important question remains: can the geometric, rearrangement-induced phase transitions described above be observed experimentally? In addressing this question, two physical constraints must be taken into account. First, once a grain has been sliced, cut, or crushed, the resulting fragments cannot spontaneously reattach to reconstitute the original solid grain. Second, the levels of a tower are not mechanically bonded to one another; instead, the upper levels merely rest on the lower ones.
@@
A tower at fragmentation stage $i$ is constructed from grains of length $l_{i}=2^{\,n-i+1}a$ and has a height $h_{i}=2^{\,i-1}a$ (see Eqs. (6) and (7)). The conjugate tower at stage ni+2n-i+2 is instead constructed from grains of length $l_{n-i+2}=2^{\,i-1}a$ and has a height $h_{n-i+2}=2^{\,n-i+1}a$. At first sight, the two towers appear to be assembled from identical grains: in the former case the grains are arranged horizontally, whereas in the latter they are aligned vertically. Crucially, however, this apparent equivalence is misleading. The vertically oriented “grain” is not a single object, but rather a stack of smaller fragments, whereas the horizontally oriented grain is an intact solid piece.
@@
For illustration, the case $n=2$, shown in Fig. 3, may be considered. At fragmentation stage $i=1$, the grain length is $l_{1}=2^{\,2-1+1}a=4a$, and the tower consists of a single level of height $h_{1}=2^{\,1-1}a=a$. At the conjugate stage ni+2=3n-i+2=3, the grain length is $l_{3}=2^{\,i-1}a=a$, and the tower height is $h_{3}=2^{\,2-1+1}a=4a$. Four unit blocks are stacked vertically, producing an object that appears identical, at a coarse-grained geometric level, to a single grain of length $4a$ at stage i=1i=1\,.
@@
It is therefore concluded that the geometric phase transition is intrinsically directional and may occur only in the direction
$i\;\longrightarrow\;n-i+2\,$,
unless the grains at stage $i$ are already arranged and oriented in the geometric configuration corresponding to stage ni+2n-i+2\,.
With this constraint in mind, the following experimental situation may be considered. Suppose that $N_{g}$ grains, each of length $l_{g}$, are given. The fragmentation stage $i$ of the hypothetical parental object corresponding to this configuration is determined by Eq. (16). The conjugate stage $i_{c}\equiv n-i+2\,$, expressed directly in terms of experimentally measurable parameters, is then given by
@@
The implications of this criterion may be illustrated by reconsidering the example of the nylon rods discussed above. In this case, the phase-transition condition implies $N_{g}<16\,$. As a further example with a larger aspect ratio, matchsticks (with tips removed), characterised by a cross-sectional side length $a=2\,\mathrm{mm}$ and a length $l_{g}=5\,\mathrm{cm}\,$, yield the condition $N_{g}<100\,$. In both cases, the number of grains required for the observation of the transition is relatively small.
@@
It is therefore suggested that such geometric phase transitions are restricted to mesoscopic systems, which may explain why they are not commonly observed in everyday granular materials.
The phase-transition criterion (21) also provides additional insight into the interpretation of Eq. (17). When this criterion is not satisfied, the system admits a unique geometric configuration, and repeated experimental measurements of the relative volume $R$ are expected to form a single cluster centered around the characteristic value prescribed by Eq. (17). If the criterion is satisfied, however, two conjugate geometric configurations become accessible. In this regime, repeated preparations of the system may populate either configuration, resulting in two distinct clusters in the measured values of $R$, each corresponding to one of the conjugate geometric phases.
@@
The observations reported in Ref. [4] support the notion that geometric phase transitions of the type discussed here may be induced by mechanical agitation. In that study, pistachio shells - whose shapes deviate significantly from the idealized grains considered in the present model - were found to interlock, nest, and compact under shaking. This behaviour indicates that such transitions are unlikely to occur uniformly throughout the sample, but instead arise locally within small clusters of grains. This naturally motivates consideration of clusters whose grain number and aspect ratio satisfy the phase-transition criterion derived above. Local rearrangements of this type may then cumulatively contribute to the macroscopic volume reduction observed during agitation. To test this hypothesis experimentally, the bulk packing density may be modelled as a superposition of contributions from different geometric configurations (or phases), with the evolution of this distribution monitored as a function of the agitation protocol.
@@
Although the present model is purely geometric and entirely static in nature, any experimental realization necessarily involves real grains subject to friction, gravity, and contact constraints. Gentle mechanical agitation is therefore introduced not to endow the system with thermodynamic character, but simply to facilitate internal rearrangements and allow the system to explore its admissible geometric configurations. In this sense, agitation serves as a practical means of reducing kinetic trapping and history dependence, rather than as a source of thermalization or energetic driving.
Experimental protocol (conceptual).
A minimal experimental test of the proposed geometric phase transition could proceed as follows. A collection of $N_{g}\gg 1$ elongated grains of uniform length $l_{g}$ and square cross section $a\times a$ (e.g., rods or trimmed matchsticks) is prepared. The grains are assembled in a loosely confined container that permits internal rearrangement under external agitation while suppressing large-scale translation of the sample as a whole. For fixed values of $N_{g}$ and $l_{g}$, the system is repeatedly prepared under identical macroscopic conditions and subjected to controlled mechanical excitation (tapping or shaking) of moderate intensity. Such excitation is intended solely to overcome frictional constraints and enable rearrangements between geometric configurations, rather than to induce thermal equilibration. After each preparation, the bulk occupied volume is measured and the corresponding relative volume $R$ is computed.
If the criterion (21) is not satisfied, the measurements are expected to yield a single reproducible cluster of $R$ values. If the criterion is satisfied, two distinct clusters of $R$ values may be observed across repeated preparations, corresponding to the paired conjugate configurations. The appearance of such bi-modality would indicate a geometric phase transition driven by configurational rearrangement rather than by changes in material properties. The separation between the two clusters should correlate with the prediction of Eq. (20).
These considerations should be taken into account when testing the dependence of the relative volume on grain size, Eq. (17). In the parameter regime where the criterion (21) is satisfied, two distinct branches are expected, corresponding to conjugate configurations. Outside this regime, where the criterion is not satisfied, a single branch should be observed.

### v2

@@ III.5 Remarks on paired configurations and “phase” interpretation
The pairing structure is closely related to the symmetry properties of the tower configurations.
@@
Since $i$ must be a positive integer, this condition requires careful interpretation. If $n$ is even, Eq. (13) yields an integer value, and the pair (i,n+2i)(i,n+2-i) reduces to $(n/2+1,n/2+1)$. In this case, two conjugate towers coincide and form a single self-conjugate tower. If $n$ is odd, such a reduction does not occur, and a self-conjugate tower never exists.
Conjugate towers are characterized by a rearrangement of identical-looking building blocks. When these towers coincide in the even-$n$ case, the resulting self-conjugate tower inherits properties of both configurations. It can therefore be inferred that the self-conjugate tower may be constructed from the same building blocks in two distinct orientations, implying the presence of an internal symmetry, in agreement with the behaviour already observed in Sec. III.3.
These paired configurations provide a simple geometric analogue of distinct phases. While they are composed of identical-looking building blocks, they differ by a global rearrangement that modifies the enclosed cavity volume. Rearrangement-induced transitions between conjugate towers thus correspond to transitions between geometrically distinct configurations with different effective densities.
The geometric transitions observed in the present model admit a natural interpretation in the language of phase transitions. The relative volume $R_{\,n}^{\,i}$ plays the role of a macroscopic order parameter distinguishing tower configurations. Transitions between conjugate towers involve discontinuous changes in $R_{\,n}^{\,i}$ and may therefore be regarded as first-order-like.
In contrast, when $n$ is even, the middle tower is self-conjugate: rearrangement of the building blocks alters their internal orientation without changing the overall geometry or volume. This transition is associated with a discrete rotational symmetry and corresponds to a change in internal symmetry rather than in macroscopic structure. In this sense, the self-conjugate tower represents a geometric analogue of a second-order phase transition.
@@
IV Connection to granular systems
To connect the present model with real granular systems, Eq. (10) is employed to evaluate the relative volume accessible to grains in experimental systems, while Eq. (12) is employed to identify the geometric phase transitions predicted by the model. Both expressions are formulated in terms of the discrete parameters $n$ and $i$. To enable experimental comparison, these parameters must be re-expressed in terms of measurable grain characteristics.
@@ IV.1 Remarks on experimental observation of the relative volume
A granular sample composed of identical grains with square cross-section of side length $a$ is considered. It is further assumed that these grains may be regarded as having been produced by sequential slicing of a hypothetical elongated parent object, in accordance with the fragmentation rule introduced earlier.
By comparing the total mass of the granular sample with the mass of an individual grain, the total number of grains in the sample, denoted by $N_{g}$, is determined. Let $l_{g}$ denote the length of an individual grain. The length $l$ of the hypothetical parent object is then given by
@@
The next step is to identify the fragmentation stage of the hypothetical parent object at which the size of a model grain coincides with that of the grains used in the experiment. At fragmentation stage $i$, the grain length $l_{i}$ (given by Eq. (6)) must match the experimental grain length $l_{g}$, so that
@@
As a consistency check, for a single grain ($N_{g}=1$), Eq. (16) yields $i=0$, corresponding to the absence of fragmentation, as expected.
Substituting Eqs. (15) and (16) into Eq. (10) yields
@@
Although the relative volume $R$ is expressed in Eq. (17) as a single function of the geometric ratio $l_{g}/a$, this quantity should be understood as a characteristic upper bound rather than as a uniquely reproducible experimental outcome. Under nominally identical preparation conditions, repeated realizations of the system are expected to yield a distribution of measured relative volumes clustered below this bound.
For a granular assembly composed of grains with square cross section, the relative volume may therefore be determined experimentally by measuring only the grain length $l_{g}$ and the side length $a$ of the cross section. Upon successive fragmentation, these measurements may be repeated until the terminal stage is reached, at which $l_{g}=a$. In this limit, the relative volume, given by Eq. (17), approaches the value $5/4$, in agreement with Eq. (11).
Exact quantitative agreement with the theoretical curve is not expected under experimental conditions. Nevertheless, the measured values are expected to follow the qualitative trend predicted by Eq. (17) and to remain bounded above by the maximal-volume curve provided by the model. Assuming square cross sections, these considerations may be summarised as follows.
@@
If a grain has length $l_{g}$ and a square cross section with side length $a$, the upper bound of the relative volume is determined solely by the ratio $l_{g}/a$ according to Eq. (17).
@@
The upper bound of the relative volume decreases monotonically as the grain length $l_{g}$ decreases. Its absolute minimum is attained when $l_{g}=a$, at which point the upper bound equals $5/4$, a value referred to as Liza’s limit or Liza’s number. The upper bound is independent of the number of grains and always exceeds unity, corresponding to the volume obtained by packing all grains adjacently without voids.
Equation (17) may be tested against existing experimental data. In Ref. [10], experiments were conducted on monodisperse nylon 6/6 cylindrical rods with diameter $1.8\,\mathrm{mm}$ and length $7.0\,\mathrm{mm}$. Approximately 7200 rods were poured into a vertical glass tube of height $1\,\mathrm{m}$ and inner diameter $1.90\,\mathrm{cm}$. Although the present model is formulated for grains with square cross section, a reasonable estimate is expected to be obtained for cylindrical rods of comparable transverse dimensions. The rods are therefore approximated as grains with square cross section of side length $a=1.8\,\mathrm{mm}$ and grain length $l_{g}=7.0\,\mathrm{mm}$. Substitution into Eq. (17) then yields
@@
Experimentally, the quantity reported is the packing fraction $\rho$ (also commonly denoted $\phi$), defined as the fraction of the available volume occupied by the particles. When the longest grain dimension is small compared to the characteristic dimensions of the container — in the present case, the rod length relative to the tube diameter — the packing fraction may be approximated as $\rho\simeq R^{-1}$. This estimate gives $\rho\simeq 0.51$, which lies within the experimentally observed range $\rho=0.49$$0.55$ reported in Ref. [10].
IV.2 Comparison with experimental scaling laws
Experimental measurements have shown that the packing fraction $\phi$ of randomly oriented elongated particles decreases with increasing aspect ratio according to an asymptotic scaling law [8, 13]
$2\phi\times\alpha=\langle\gamma\rangle,\qquad\alpha\gg 1,$ (18)
where $\alpha=L/D$ is the aspect ratio, $L$ is the rod length, $D$ is the rod diameter, and $\langle\gamma\rangle$ is the average number of contacts per particle. Experiments indicate $\langle\gamma\rangle\approx 10.8$ for randomly packed rods [13].
The present geometric model can be compared with this scaling in the limit of large aspect ratio. For a grain with length $l_{g}$ and square cross-section of side $a$, the aspect ratio is $\alpha=l_{g}/a$. The packing fraction is related to the relative volume by $\phi=1/R$. Substituting Eq. (17) yields
$\phi=\frac{1}{R}=\frac{1}{1+l_{g}/(4a)}=\frac{4}{4+\alpha}.$ (19)
For large aspect ratios ($\alpha\gg 1$), this reduces to
$\phi\approx\frac{4}{\alpha},\qquad\text{or equivalently}\qquad 2\phi\times\alpha\approx 8.$ (20)
This result demonstrates that the present purely geometric model captures the correct $\phi\propto 1/\alpha$ scaling observed experimentally. The numerical coefficient differs from the experimental value ($8$ versus $\approx 10.8$), which is expected given that the model provides an upper bound on accessible volume and assumes maximal-volume configurations rather than typical random packings. The agreement in scaling behavior confirms that volume evolution under fragmentation, as described by the present model, is consistent with the experimental observation that packing densities are size-invariant and have a purely geometric origin [13].
Figure 5 compares the model predictions with experimental data from Freeman et al. [9] for aspect ratios ranging from $\alpha=4$ to $\alpha=32$. The experimental packing fractions lie consistently above the theoretical curve, as expected for a model that provides an upper bound on relative volume (or equivalently, a lower bound on packing fraction). The agreement in scaling behavior ($\phi\propto 1/\alpha$) validates the geometric approach.
$\alpha$$\phi$05101520253000.10.20.30.40.50.6Model: $\phi=4/(4+\alpha)$Experimental $\phi_{\infty}$Asymptotic: Eq. (18)
Figure 5: Experimental packing fraction $\phi_{\infty}$ as a function of aspect ratio $\alpha$ (circles with downward error bars) from Freeman et al. [9] compared with the present model (Eq. (19), solid blue line) and the asymptotic scaling law (Eq. (18), dashed blue line). The model provides a lower bound on packing fraction, while the asymptotic scaling law provides an empirical upper bound based on contact statistics.
The asymptotic scaling law (Eq. (18)), shown as a dashed line in Fig. 5, provides an empirical upper bound on packing fraction based on contact statistics in random packings. The experimental data lie between the two theoretical curves: above the geometric lower bound from the present model and below (or near) the empirical upper bound from the asymptotic scaling law. This positioning is physically reasonable, as real random packings achieve packing fractions intermediate between the maximal-volume geometric configurations considered here and the contact-limited configurations reflected in the asymptotic scaling. The convergence of both theoretical curves at large $\alpha$, with experimental data consistently confined between them, confirms that geometric packing constraints dominate throughout the range of aspect ratios studied, with the asymptotic regime exhibiting universal scaling behavior.
IV.3 Remarks on experimental observation of geometric phase transitions
@@
An important question remains: can the geometric, rearrangement-induced phase transitions described above be observed experimentally? In addressing this question, two physical constraints must be taken into account. First, once a grain has been sliced, cut, or crushed, the resulting fragments cannot spontaneously reattach to reconstitute the original solid grain. Second, the levels of a tower are not mechanically bonded to one another; the upper levels merely rest on the lower ones under gravity.
A tower at fragmentation stage $i$ is constructed from grains of length $l_{i}=2^{\,n-i+1}a$ and has a height $h_{i}=2^{\,i-1}a$ (see Eqs. (6) and (7)). The conjugate tower at stage ni+2n-i+2 is instead constructed from grains of length $l_{n-i+2}=2^{\,i-1}a$ and has a height $h_{n-i+2}=2^{\,n-i+1}a$.
@@
At first glance, the two towers appear to be assembled from identical grains: in the former case the grains are arranged horizontally, whereas in the latter they are aligned vertically. Crucially, however, this apparent equivalence is misleading. The vertically oriented “grain” is not a single object, but rather a stack of smaller fragments, whereas the horizontally oriented grain is an intact solid piece.
@@
For illustration, the case $n=2$, shown in Fig. 3, may be considered. At fragmentation stage $i=1$, the grain length is $l_{1}=2^{\,2-1+1}a=4a$, and the tower consists of a single level of height $h_{1}=2^{\,1-1}a=a$. At the conjugate stage ni+2=3n-i+2=3, the grain length is $l_{3}=2^{\,i-1}a=a$, and the tower height is $h_{3}=2^{\,2-1+1}a=4a$. Four unit blocks are stacked vertically, producing an object that appears identical, at a coarse-grained geometric level, to a single grain of length $4a$ at stage $i=1$.
@@
It is therefore concluded that the geometric phase transition is intrinsically directional and may occur only in the direction $i\to n-i+2$. However, if the intact grains at stage $i$ happen to be oriented vertically — matching the orientation of the stacked fragments at stage ni+2n-i+2 — then rotation to horizontal alignment enables the reverse transition $n-i+2\to i$ without requiring recombination of fragments.
With this constraint in mind, the following experimental situation may be considered. Suppose that $N_{g}$ grains, each of length $l_{g}$, are given. The fragmentation stage $i$ of the hypothetical parent object corresponding to this configuration is determined by Eq. (16). The conjugate stage $i_{c}\equiv n-i+2$, expressed directly in terms of experimentally measurable parameters, is then given by
@@
The phase-transition criterion (24) provides insight into the interpretation of Eq. (17). When this criterion is not satisfied, the system admits a unique geometric configuration, and repeated experimental measurements of the relative volume $R$ are expected to form a single cluster centred around the characteristic value prescribed by Eq. (17). If the criterion is satisfied, however, two conjugate geometric configurations become accessible. In this regime, repeated preparations of the system may populate either configuration, resulting in two distinct clusters in the measured values of $R$, each corresponding to one of the conjugate geometric phases.
@@
The implications of this criterion may be illustrated by concrete examples. For the nylon rods discussed in Sec. IV.1 (with $l_{g}=7.0\,\mathrm{mm}$ and $a=1.8\,\mathrm{mm}$), the phase-transition condition implies $N_{g}<16$. As a further example with larger aspect ratio, matchsticks with tips removed, characterised by a cross-sectional side length $a=2\,\mathrm{mm}$ and length $l_{g}=5\,\mathrm{cm}$, yield the condition $N_{g}<100$.
@@
In both cases, the number of grains required for the observation of the transition is relatively small. It is therefore concluded that geometric phase transitions of the type described here are restricted to mesoscopic systems, which may explain why they are not commonly observed in everyday granular materials.
Experimental support for the local nature of these transitions may be found in X-ray tomography studies. Freeman et al. [9] observed local alignment of small groups of rods (forming what may be termed domains) in containers filled with grains of aspect ratio $\alpha=8$. For these rods in containers of diameter $\approx 4.9\,l_{g}$, the criterion (24) requires $N_{g}<32$ for geometric phase transitions to occur. The observed local domains, while themselves randomly oriented within the container, are consistent with geometric rearrangements occurring in small grain clusters rather than uniformly throughout the sample. Similarly, Zakine et al. [5] found that pistachio shells — whose shapes deviate significantly from the idealised grains considered here — compact and interlock under mechanical agitation, suggesting that such transitions are facilitated by local rearrangements within small clusters.
@@
This naturally motivates consideration of domains whose grain number and aspect ratio satisfy the phase-transition criterion derived above. Local rearrangements of this type may then cumulatively contribute to the macroscopic volume reduction observed during agitation. To test this hypothesis experimentally, the bulk packing density may be modelled as a superposition of contributions from different geometric configurations (or phases).
@@
Although the present model is purely geometric and entirely static in nature, any experimental realisation necessarily involves real grains subject to friction, gravity, and contact constraints. Gentle mechanical agitation is therefore introduced not to endow the system with thermodynamic character, but simply to facilitate internal rearrangements and allow the system to explore its admissible geometric configurations. In this sense, agitation serves as a practical means of reducing kinetic trapping and history dependence, rather than as a source of thermalisation or energetic driving.
Experimental signatures.
The geometric phase transition mechanism admits two distinct experimental regimes depending on the total grain number.
Small-system regime ($N_{g}\lesssim 4l_{g}/a$): When the criterion (24) is satisfied for the entire grain assembly, repeated preparations under controlled agitation should yield a bimodal distribution of bulk relative volumes $R$, with peaks separated by $\Delta R$ (Eq. (23)) corresponding to global conjugate configurations.
Large-system regime ($N_{g}\gg 4l_{g}/a$): When the total grain number exceeds the criterion, global phase transitions are prohibited, but local transitions within domains containing $N_{g}\lesssim 4l_{g}/a$ grains remain possible. In this regime, the bulk relative volume forms a single distribution, but X-ray tomography should reveal aligned grain clusters (domains) of characteristic size $N_{\text{domain}}\sim 4l_{g}/a$. Since this constraint scales linearly with the aspect ratio $\alpha=l_{g}/a$, experiments conducted on grains of systematically varying aspect ratios should reveal that characteristic domain sizes increase proportionally with $\alpha$. Observation of such scaling would provide direct experimental evidence for the geometric phase transition mechanism, while its absence would indicate alternative origins for the local alignment patterns.